# RoboFusion: Towards Robust Multi-Modal 3D Object Detection via SAM Appendix A Appendix A.1 Broader Impacts

### v2

@@ A.1 Broader Impacts
Our work aims to develop a robust framework to address out-of-distribution (OOD) noise scenarios in autonomous driving (AD). To the best of our knowledge, RoboFusion is the first method that leverages the generalization capabilities of visual foundation models (VFMs) like SAM sam, FastSAM fastsam, and MobileSAM mobilesam for multi-modal 3D object detection. Although existing multi-model 3D object detection methods achieve the state-of-the-art (SOTA) performance of ‘clean’ datasets, they overlook the robustness of real-world scenarios. Therefore, we believe it is valuable to combine VFMs and multi-modal 3D object detection to mitigate the impact of OOD noise scenarios.
@@ A.2 More Results
A.2.1 Roles of Different Modules in RoboFusion.
To assess the roles of different modules in RoboFusion, we conduct an ablation study on the original SAM rather than SAM-AD, as shown in Table 1, where a) is the results of the baseline focalconv, b)-e) shows the performance of our RoboFusion-L under different modules. According to Table 1, SAM and AD-FPN modules significantly improve the performance in OOD noisy scenarios. It is worth noticing that DGWA module significantly improves the performance, especially in snow noisy scenarios. By Table 1, the impact of fog noise on point clouds is relatively minor. But, using A.F. (Adaptive Fusion) module to dynamically aggregate point cloud features and image features exhibits significant enhancements in fog-noise scenarios.
@@
†: Results from Ref. Robustness3d.
@@
†: Results from Ref. Robustness3d.
@@
†: Results from Ref. Robustness3d.
@@
A.2.2 More Results on the KITTI-C validation set.
Besides the experimental results mentioned in the main text, we follow Ref. Robustness3d to extend our work to a wider range of noise scenarios, including Gaussian, Uniform, Impulse, Moving Object, Motion Blur, Local Density, Local Cutout, Local Gaussian, Local Uniform, and Local Impulse, as shown in Tables 2, 3, and 4. From these Tables, compared with multi-modal methods including EPNet epnet, Focals Conv focalconv,and LoGoNet logonet, our RoboFusion-L, RoboFusion-B, and RoboFusion-T present better results across the most of noise scenarios, perform the best on average.
In addition, to provide a clearer analysis of performance across different noise severities, we present a more detailed comparative study of our RoboFusion-L and LoGoNet on the KITTI-C dataset, as shown in Table 5. It is evident that our method demonstrates a high degree of robustness, exhibiting the most stable results with the variance of noise severities. For instance, when considering snow conditions, the performance of our RoboFusion-L shows a marginal variation from 86.69% to 83.67% across severities from 1 to 5. In contrast, LoGoNet’s performance drops from 55.07% to 45.02% over the same severity range. Furthermore, in the presence of moving object noise, our method outperforms LoGoNet. In summary, our RoboFusion exhibits remarkable robustness and generalization capabilities, making it well-suited to diverse noise scenarios.
@@
†: Results from Ref. Robustness3d.
@@
A.2.3 More Results on the nuScenes-C validation set.
As depicted in Table 6, compared to SOTA methods, including PointPillars pointpillars, CenterPoint centerpoint, FCOS3D fcos3d, DETR3D detr3d, BEVFormer li2022bevformer, FUTR3D chen2023futr3d, TransFusion transfusion, BEVFusion bevfusion-mit, and DeepInteraction deepinteraction, our RoboFusion demonstrates superior performance across 20 noise scenarios in AD on average. Overall, our method exhibits not only exceptional robustness in weather-induced noise scenarios, but also shows remarkable resilience across a broader noise include sensor, motion and object noise.
@@ A.3 Visualization
As shown in Fig. 1, we provide visualization results between our RoboFusion-L and LoGoNet on the KITTI-C dataset. Overall, compared to SOTA methods like LoGoNet logonet, our method enhances the robustness of multi-modal 3D object detection by leveraging the generalization capability and robustness of VFMs to mitigate OOD noisy scenarios in AD.
@@ A.4 More Limitations
Although we have mentioned the two main limitations in the ‘Conclusions’ section of the main text, our RoboFusion still has other limitations. Our method does not achieve the best performance in all noisy scenarios. For instance, as shown in Table 2, our method does not show the best in ‘Moving Object’ noisy scenarios. Furthermore, we conduct experiments only on the corruption datasets Robustness3d rather than real-world datasets. It is valuable to construct a real-world corruption dataset, but it must be an expensive work.

### v4

@@ A.1 Broader Impacts
Our work aims to develop a robust framework to address out-of-distribution (OOD) noise scenarios in autonomous driving (AD). To the best of our knowledge, RoboFusion is the first method that leverages the generalization capabilities of visual foundation models (VFMs) like SAM Kirillov et al. (2023), FastSAM Zhao et al. (2023), and MobileSAM Zhang et al. (2023a) for multi-modal 3D object detection. Although existing multi-model 3D object detection methods achieve the state-of-the-art (SOTA) performance of ‘clean’ datasets, they overlook the robustness of real-world scenariosSong et al. (2024a). Therefore, we believe it is valuable to combine VFMs and multi-modal 3D object detection to mitigate the impact of OOD noise scenarios.
@@ A.2 More Results
A.2.1 Specific classes AP on the nuScenes-C validation set.
As shown in Table 9, we present a comparison of Specific classes AP between TransFusion and our RoboFusion-L on the nuScenes-C validation set, encompassing scenarios with snow, rain, fog, and strong sunlight noise. It is evident from the results that RoboFusion-L exhibits superior performance.
| mAP | Car | Truck | C.V. | Bus | Trailer | Barrier | Motor. | Bike | Ped. | T.C. | ||
| Snow | T.F. | 63.30 | 84.55 | 58.41 | 25.50 | 62.31 | 56.00 | 70.19 | 69.98 | 43.69 | 84.24 | 78.16 |
| R.F. | 67.12 | 87.21 | 60.88 | 29.47 | 67.45 | 58.99 | 75.12 | 71.45 | 48.28 | 86.23 | 86.12 | |
| +3.82 | +5.14 | +4.93 | +4.59 | +7.96 | ||||||||
| Rain | T.F. | 63.35 | 85.37 | 56.87 | 25.12 | 64.65 | 55.10 | 71.99 | 68.21 | 44.13 | 83.87 | 78.14 |
| R.F. | 67.58 | 86.79 | 60.44 | 30.21 | 65.41 | 58.12 | 75.47 | 71.39 | 50.87 | 88.91 | 88.23 | |
| +4.23 | +5.09 | +6.74 | +5.04 | +10.09 | ||||||||
| Fog | T.F. | 53.67 | 80.23 | 48.51 | 18.04 | 50.69 | 53.03 | 62.24 | 54.53 | 25.27 | 80.63 | 66.48 |
| R.F. | 67.01 | 87.56 | 59.03 | 29.36 | 66.10 | 57.23 | 74.33 | 72.01 | 49.50 | 87.08 | 87.91 | |
| +13.34 | +10.52 | +11.32 | +15.41 | +12.09 | +17.48 | +24.23 | +21.43 | |||||
| S.L. | T.F. | 55.14 | 81.99 | 48.07 | 19.78 | 51.09 | 52.57 | 63.68 | 55.09 | 26.98 | 82.68 | 69.49 |
| R.F. | 67.24 | 87.67 | 57.74 | 31.00 | 64.29 | 58.94 | 75.23 | 70.23 | 50.82 | 88.70 | 87.82 | |
| +12.10 | +11.22 | +13.20 | +11.55 | +15.14 | +23.30 | +18.33 |
A.2.2 Roles of Different Modules in RoboFusion.
To assess the roles of different modules in RoboFusion, we conduct an ablation study on the original SAM rather than SAM-AD, as shown in Table 10, where a) is the results of the baseline Chen et al. (2022), b)-e) shows the performance of our RoboFusion-L under different modules. According to Table 10, SAM and AD-FPN modules significantly improve the performance in OOD noisy scenarios. It is worth noticing that DGWA module significantly improves the performance, especially in snow noisy scenarios. By Table 10, the impact of fog noise on point clouds is relatively minor. But, using A.F. (Adaptive Fusion) module to dynamically aggregate point cloud features and image features exhibits significant enhancements in fog-noise scenarios.
@@
†: Results from Ref. Dong et al. (2023).
@@
†: Results from Ref. Dong et al. (2023).
@@
†: Results from Ref. Dong et al. (2023).
@@
A.2.3 More Results on the KITTI-C validation set.
Besides the experimental results mentioned in the main text, we test our RoboFusion on KITTI-C and nuScenes-C Dong et al. (2023) to extend our work to a wider range of noise scenarios, including Gaussian, Uniform, Impulse, Moving Object, Motion Blur, Local Density, Local Cutout, Local Gaussian, Local Uniform, and Local Impulse, as shown in Tables 11, 12, and 13. From these Tables, compared with LiDAR-only methods including SECOND Yan et al. (2018), PointPillars Lang et al. (2019), PointRCNN Shi et al. (2019) and PV-RCNN Shi et al. (2020), Camera-Only methods including Smoke Liu et al. (2020b), ImVoxelNet Rukhovich et al. (2022), and multi-modal methods including EPNet Huang et al. (2020), Focals Conv Chen et al. (2022), and LoGoNet Li et al. (2023), our RoboFusion-L, RoboFusion-B, and RoboFusion-T consistently outperform across various noise scenarios and achieve the best overall performance. Overall, our RoboFusion demonstrates superior performance in weather-noisy (i.e. Snow, Rain, Fog, and Strong Sunlight) scenarios and exhibits better results across a broader range of scenarios, which shows remarkable robustness and generalizability.
A.2.4 Performance Comparison Analysis with the LoGoNet.
In addition, to provide a clearer analysis of performance across different noise scenarios, we present a more detailed comparative study of our RoboFusion-L and LoGoNet Li et al. (2023) on the KITTI-C validation dataset, as shown in Table 14. It is worth noting that LoGoNet is a SOTA multi-modal 3D detector known for its exceptional robustness and high accuracy. Dong et al. (2023) provides noise at varying levels, with the KITTI-C dataset including 5 severities. It is evident that our method demonstrates a high degree of robustness, exhibiting the most stable results with the variance of noise severities. For instance, when considering snow conditions, the performance of our RoboFusion-L shows a marginal variation from 86.69% to 83.67% across severities from 1 to 5. In contrast, LoGoNet’s performance drops from 55.07% to 45.02% over the same severity range. Furthermore, in the presence of moving object noise, our method outperforms LoGoNet. In summary, our RoboFusion exhibits remarkable robustness and generalization capabilities, making it well-suited to diverse noise scenarios.
@@
†: Results from Ref. Dong et al. (2023).
@@
A.2.5 More Results on the nuScenes-C validation set.
As depicted in Table 15, compared with LiDAR-only methods including PointPillars Lang et al. (2019), and CenterPoint Yin et al. (2021), Camera-Only methods FCOS3D Wang et al. (2021), DETR3D Wang et al. (2022b), and BEVFormer Li et al. (2022c) and multi-modal methods including FUTR3D Chen et al. (2023), TransFusion Bai et al. (2022), BEVFusion Liu et al. (2023) and DeepInteraction Yang et al. (2022), our RoboFusion demonstrates superior performance across more noise scenarios in AD on average. For instance, our RoboFusion-L excels in 10 noise scenarios, including Weather (Snow, Rain, Fog, Strong Sunlight), Sensor (Density, Cutout, Crosstalk), Motion (Compensation, Motion Blur), and Object (Local Cutout), outperforming DeepInteraction Yang et al. (2022) which achieves the best performance only in 5 of these noise scenarios. Overall, our method exhibits not only exceptional robustness in weather-induced noise scenarios, but also shows remarkable resilience across a broader noise include sensor, motion and object noise.
@@ A.3 Visualization
As shown in Fig. 5, we provide visualization results between our RoboFusion-L and LoGoNet on the KITTI-C dataset. Overall, compared to SOTA methods like LoGoNet Li et al. (2023), our method enhances the robustness of multi-modal 3D object detection by leveraging the generalization capability and robustness of VFMs to mitigate OOD noisy scenarios in AD.
@@ A.4 More Limitations
Although we have mentioned the two main limitations in the ‘Conclusions’ section of the main text, our RoboFusion still has other limitations. Our method does not achieve the best performance in all noisy scenarios. For instance, as shown in Table 11, our method does not show the best in ‘Moving Object’ noisy scenarios. Furthermore, we conduct experiments only on the corruption datasets Dong et al. (2023) rather than real-world datasets. It is valuable to construct a real-world corruption dataset, but it must be an expensive work.